the shortest wavelength limit shifts by 26 pm when the operating voltage in an X- RAY TUBE IS INCREASED TO 1.5 TIMES THE original value. the initial value of operating voltage.

Grade:

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
lambda_min(Angs)=12375/V
26*10^(-2)=12375(1/V-1/1.5V)
solve for V
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty